40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16266     Entry time: Thu Jul 29 14:51:39 2021
Author: Paco 
Type: Update 
Category: Optical Levers 
Subject: Recenter OpLevs 

[yehonathan, anchal, paco]

Yesterday around 9:30 pm, we centered the BS, ITMY, ETMY, ITMX and ETMX oplevs (in that order) in their respective QPDs by turning the last mirror before the QPDs. We did this after running the ASS dither for the XARM/YARM configurations to use as the alignment reference. We did this in preparation for PRFPMI lock acquisition which we had to stop due to an earthquake around midnight

ELOG V3.1.3-